# Quantum Corrections to Symmetron Fifth-Force Profiles

### v1

@@ 5 Tadpole Contribution and Quantum Corrections
We are now able to compute the tadpole contribution, which is essentially the renormalised coincidence limit of the Green’s function we found in the last section. This time, we do need to integrate over the modes from the continuum approximation, resulting in a divergent loop integral that we regularise with a simple momentum cutoff. While the counterterms can be found exactly, the renormalised tadpole contribution can only be found via a numerical method, by subtracting divergences at the level of the integrand in the loop integral [40, 42].
@@ 5.1 Coincidence Limit
Following from (4.78), the coincidence limit of the Green’s function in the planar-wall approximation takes the form
@@
where $G(\mathbf{x};E)\equiv G(\mathbf{x},\mathbf{x};E)$ and $G(s;E,k)\equiv G(s,s;E,k)$. Recall that the $R^{2}$ prefactor follows from adjustments made when we took the thin-wall approximation (see eq. (4.10)). Adding back time dependence and taking the coincident limit in time yields
@@
where $G(x)\equiv G(x,x)$, and we have exploited the fact that $G(s;E,k)$ is an even function in $E$. This loop integral is well defined after taking a Wick rotation (at least exterior to the source), the action of which is equivalent to replacing every occurrence of $E^{2}$ with $E^{2}$ at the cost of a factor of $i$,
@@
Figure 4: The coincident limit of the Green’s function, where we have chosen $p=\gamma$. We have set $\lambda=0.1$ and $\rho/\mu^{2}M^{2}=25$.
@@
To evaluate eq. (5.4), we first analyse the exterior part, for which we use a subscript ++. Before Wick rotating, it is
@@
We split this expression into a part $G_{+}^{(2)}$ which explicitly depends on the value of the field at the surface of the source and a part $G_{+}^{(1)}$ which does not (at least when taken as a function of the field itself). The latter may be treated exactly and is given by
@@
This is exactly $i$ times the result obtained from the Euclidean calculation [40], as we might expect given the similarity of the external part of the classical solution and the resulting differential problem to that of determining the quantum corrected bounce configurations in the Coleman description of thin-wall vacuum decay. The second term $G_{+}^{(2)}$ is too complicated to be amenable to a similar analysis. However, the loop integral of $G_{+}^{(2)}$ converges, and thus all we need to know is that it will contribute some finite shift which can be calculated using numerical methods.
@@
Upon doing so, the term which does not contain any unevaluated Wrońskians is
@@
Note the singularities: simple poles at $E=\pm k$ and $E=\pm\sqrt{k^{2}-3g^{2}\gamma^{2}}$ and branch points at $E=\pm\sqrt{k^{2}+g^{2}\gamma^{2}}$. Note further that, for all $k$, the pair of branch points and first pair of poles lie on the real energy axis, whereas the second pair of poles lie on the imaginary energy axis if $k^{2}<3g^{2}\gamma^{2}$. In this case, the $i\epsilon$ prescription deflects the real singularities as before (positive below and negative above the real line), but the imaginary poles undergo an infinitesimal anticlockwise rotation, as shown in Figure 6. We seem to have a propagator whose mass takes the wrong sign, i.e., a tachyon, which contributes a residue with an additional factor of $i$. This tachyonic mode turns out to be spurious — a fact corroborated by our considerations in Appendix A and continuity with the exterior contribution — but prevents us from calculating $G_{-}^{(2)}$ in closed form, at least by any method which we are aware of. Consequently, renormalisation will have to be performed numerically. We elaborate on this in the next section.
@@ 5.2 Renormalisation
The one-loop contribution to the equation of motion for $\varphi$ is the first derivative of the Coleman–Weinberg effective potential [57], given by
$V_{\text{CW}}(\phi)=V(\phi)+\frac{1}{2}\delta m^{2}\phi^{2}+\frac{1}{4}\delta\lambda\phi^{4}-\frac{i}{2}\int\,\frac{\text{d}^{4}k}{\left(2\pi\right)^{4}}\,\ln\left(1-\frac{1}{k^{2}}\frac{\text{d}^{2}V}{\text{d}\phi^{2}}\right)\>,$ (5.10)
@@
$\int\,\frac{\text{d}^{4}k}{\left(2\pi\right)^{4}}\,\ln\left(1-\frac{1}{k^{2}}\frac{\text{d}^{2}V}{\text{d}\phi^{2}}\right)$ $=\frac{i}{2\pi^{2}}\int_{0}^{\Lambda}\text{d}p\,p^{2}\left(\sqrt{p^{2}+\frac{\text{d}^{2}V}{\text{d}\phi^{2}}}-p\right)\>.$
Imposing the following renormalisation conditions [40],
@@
Since we are unable to determine the UV-divergent part of the interior Green’s function in closed form, we will need to integrate the Green’s function numerically — using standard functions in Mathematica. To effect the renormalisation, we introduce pseudo-counterterms that allow us to subtract the divergences at the level of the integrand [40]. These are functions $\Delta m^{2}(p)$ and $\Delta\lambda(p)$ which satisfy the integral equations
@@
since it is terms like these which give rise to logarithmic divergences, but not quadratic divergences. Integrating and matching like terms fixes the constants to be
@@
Figure 7: The relative error $\eta$ between the exact and numerical calculations of the integral of $G_{+}^{(1)}$. Note that the horizontal axis begins at $\phi_{0}/v$.
@@ 5.3 One-Loop Correction
With the tadpole contribution successfully renormalised, we can now solve the equation of motion for the quantum correction $\delta\phi$. The renormalised correction to the classical field configuration $\delta\phi$ is time dependent in general [(2.23)] and satisfies
@@
Notice that the differential operator is exactly that which defines the Green’s function for $k=E$, equivalently $n=2$. Therefore, this equation is solved by the convolution
@@
For the parameters used to generate Figure 8, this amounts to a shift of about $8.5\%$ in the vacuum expectation value, which is precisely what we observe in Figure 9. The more subtle feature hiding in Figure 8 is that the quantum field is not simply a rescaled version of the classical field. The derivative of the field profile, when normalised relative to its VEV, has also shifted. Figure 10 shows this. The mass roughly gives the rate with which the field interpolates between its two asymptotes. Thus, it follows that the shift in the derivative is due to the shift in the relationship between the Lagrangian mass parameter $\mu$ and the physical mass $m$. At tree level, the physical mass is just the previously derived mass of scalar fluctuations,
@@
Since all functions here are spherically symmetric, it is sufficient to consider only the radial component of the force hereafter. We do not attempt to determine the exact way in which the mass and VEV shifts contribute to $\delta\phi$. However, from (5.33), one might naively assume that the VEV shift contributes “twice”, in a sense (since the field is squared), while the mass shift (which applies because of the derivative) contributes “once”. This very rough argument implies that the relative shift in the force or acceleration $\Delta F$ is of order
@@
From a naive perturbative expansion, one might expect an expression like $6\lambda/\pi^{2}$, which is remarkably close. This expression suggests that the shift in the force due to the self-interaction is, to a good approximation, linear in the coupling — at least in the perturbative regime — despite the non-perturbative dependence of the VEV on the coupling. Figure 13 corroborates this claim. For the parameters used to generate Figure 8, the naive calculation predicts a $30\%$ shift in the strength of the force, which is precisely what we observe, around a distance of one Compton wavelength, in Figure 12(a). This approximation holds well throughout the parameter space; the values used to generate Figure 12(b) would imply a correction of about $20\%$, which the plot agrees with. Naturally, the approximation breaks down very close to or far from the surface of the source, where one would expect quantum effects to become less relevant. Notice, however, that for large masses we observe strong deviations from tree-level even close to the source.
The quantum-corrected force is generally weaker than its classical counterpart, except after about four Compton wavelengths from the surface of the source, where the quantum force becomes stronger. While this strengthening represents a relatively large fractional change, in absolute terms, the difference is negligible. As one would expect, the shift in the force vanishes as $\lambda\rightarrow 0$. Note, however, that for sufficiently small $\lambda$ quantum corrections due to interactions with Standard Model fields are expected to become the dominant contribution, which we leave for future work.
@@
It is worth considering how the magnitude of the correction to the force depends on the other free parameters. Since changes in the mass scale $M$ are degenerate with changes in the density $\rho_{0}$ of the source, the only remaining free parameter of interest is the field mass $\mu$. Our rough arguments from earlier in this subsection would suggest that changing the mass parameter should have next to no effect on the force. This is approximately true. We observe a slight negative correlation between $\mu$ and $\Delta F$, but one that amounts to a change of about $0.4\%$ in the magnitude of the relative correction to the force for all $\mu\in(0,1]\,$GeV (the widest mass range where our numerical methods yield trustworthy results) in Figure 13(c), with similar shown in Figure 13(d). Note that we have extrapolated the linear relationship, ignoring the apparent divergence for small masses. This divergence may be a numerical artefact, a consequence of underflow errors. Alternatively, this could be a signal of the breakdown of the thin-wall approximation, since for $\mu\rightarrow 0$ we have $1/\mu\gg R$.
@@
Although these corrections seem very large in places, the strength of the implications for experiments is not necessarily so. For very light symmetron fields, not only are the forces very small, but the distance where we predict the most significant disparity between the tree-level and one-loop forces is far outside the range of tabletop experiments designed to detect them. Our results have the strongest implications for experiments which probe heavier symmetron fields.
@@ 6 Conclusions
We set out to compute the leading-order quantum corrections to the symmetron field around extended sources. To this end, we have solved for the exact field profile around spherical sources whose radii are much larger than the Compton wavelength of the field, meaning our results are also relevant for systems with planar geometry [26, 41, 31, 30, 29, 28, 27]. We have computed the inverse of the fluctuation operator (the propagator) and the renormalised tadpole contribution. With that, we derived and numerically solved the equation of motion satisfied by the one-loop correction to the classical field profile. We have found that the action of quantum corrections is to flatten the gradients of the classical field profiles and reduce their vacuum expectation values, which agrees with similar analyses of the impact of quantum corrections on tunnelling configurations [40]. As a result, forces derived from one-loop field profiles are generally weaker than those computed from tree-level profiles, for the same input parameters.
@@
In certain areas of parameter space, the classical field remains virtually identical to the quantum field. This is true in particular for very small self-couplings. One should note, however, that for $\lambda\lesssim 10^{-40}$, it is expected that couplings to the Standard Model dominate, and thus a new source of quantum corrections would have to be taken into account. For sufficiently high but still perturbative couplings, we observe that the quantum-corrected force can be considerably weaker than the classical prediction. Indeed, for regions of parameter space that sit within current constraints [53, 21, 22], the symmetron force is as much as 30% weaker than classically predicted.
@@
While our results speak mostly to perturbative self-couplings, there is a clear trend which shows that the strength of the correction scales almost linearly with the strength of the self-interaction. We suspect that this trend continues into the non-perturbative regime, and may even amount to a symmetron force that is practically undetectable for sufficiently high self-couplings. Phenomenologically, it would also be difficult to use such a field to account for observations related to dark matter. Our results also apply directly to systems with large, screened and spherical sources or systems with planar geometry. However, such an idealisation does not mean we should expect quantum corrections to be minor for systems with, say, cylindrical sources [58], or relatively small sources. Field profiles still flatten away from the thin-wall approximation [43].
@@
We emphasise two different conclusions to draw from our work. The first is that, due to renormalisation, the relationship between Lagrangian parameters and physical observables shifts from the tree-level prescription. This conclusion alone would imply that current constraints apply to slightly different regions of parameter space. The second conclusion is that the spatial variation of the fifth force also changes, as indicated by Figure 12, generally growing more slowly and peaking at a different point in space in the one-loop approximation. This could impact how we optimise geometries of future tabletop experiments [41]. The spatial variation of the quantum correction is a crucial point. Since $\delta\phi$ changes with position, there is no point in space we could choose to define the physical mass and self-coupling such that the quantum correction vanishes everywhere. Put differently, there does not exist a convenient choice of renormalisation scheme which makes the quantum corrections disappear. They cannot be fine-tuned away, as it were. The quantum fluctuations about the classical background are an intrinsic and unavoidable feature of the theory.
@@
We propose several directions for future work. For instance, our method can likely produce estimates of quantum corrections for symmetron-like models, such as the complex symmetron [59] or generalised symmetron [60]. Furthermore, it has been shown that one can get around some experimental constraints by choosing models in which screening arises at the one-loop level [61], provided the background field is constant. The methods developed in this paper would allow us to revisit this result for nontrivial backgrounds. In addition, a future investigation could consider the effect of quantum corrections for a broader range of observables, including shifts in frequency spectra [37], properties of white dwarfs (mass, radius, luminosity, etc.) [59] and gravitational lensing [60]. Finally, we hope that this or a similar analysis could be used to further explore the quantum nature of fifth force theories in general.

### v2

@@ 5 Tadpole contribution and quantum corrections
We are now able to compute the tadpole contribution, which is essentially the renormalised coincidence limit of the Green’s function that we found in the last section. To do so, we need to integrate over the modes from the continuum approximation, resulting in a divergent loop integral that we regularise with a simple momentum cutoff. While the counterterms can be found exactly, the renormalised tadpole contribution can only be found via a numerical method, by subtracting divergences at the level of the integrand in the loop integral [36, 37].
@@ 5.1 Coincidence limit
Following from eq. (4.17), the coincidence limit of the Green’s function in the planar-wall approximation takes the form
@@
where $G(\mathbf{x};E)\equiv G(\mathbf{x},\mathbf{x};E)$ and $G(s;E,k)\equiv G(s,s;E,k)$. Recall that the $R^{2}$ prefactor follows from taking the thin-wall approximation (see eq. (4.10)). Adding back the time dependence and taking the coincidence limit in time yields
@@
where $G(x)\equiv G(x,x)$, and we have exploited the fact that $G(s;E,k)$ is an even function in $E$. This loop integral is well defined after making a Wick rotation (at least exterior to the source), the action of which is equivalent to replacing every occurrence of $E^{2}$ with $E^{2}$ at the cost of a factor of $i$, giving
@@
Figure 3: The coincidence limit of the Green’s function, where we have chosen $p=\gamma$. We have set $\lambda=0.1$ and $\rho/\mu^{2}M^{2}=25$.
@@
To evaluate eq. (5.4), we first analyse the exterior part, for which we again use a subscript ++. Before Wick rotating, it is
@@
We split this expression into a part $G_{+}^{(2)}$, which explicitly depends on the value of the field at the surface of the source, and a part $G_{+}^{(1)}$, which does not (at least when taken as a function of the field itself). The latter may be treated exactly and is given by
@@
This is exactly $i$ times the result obtained from the Euclidean calculation [36], as we might expect given the similarity of the external part of the classical solution and the resulting differential problem to that of determining the quantum corrected bounce configurations in the Coleman description of thin-wall vacuum decay. The second term $G_{+}^{(2)}$ is too complicated to be amenable to a similar analysis. However, the loop integral of $G_{+}^{(2)}$ converges, and thus all we need to know is that it will contribute some finite shift, which can be calculated using numerical methods.
@@
Upon doing so, the term which does not contain Wrońskians of Legendre and hypergeometric functions is
@@
Note the singularities: simple poles at $E=\pm k$ and $E=\pm\sqrt{k^{2}-3g^{2}\gamma^{2}}$ and branch points at $E=\pm\sqrt{k^{2}+g^{2}\gamma^{2}}$. Note further that, for all $k$, the pair of branch points and first pair of poles lie on the real energy axis, whereas the second pair of poles lie on the imaginary energy axis if $k^{2}<3g^{2}\gamma^{2}$. In this case, the $i\epsilon$ prescription deflects the real singularities as before (positive below and negative above the real line), but the imaginary poles undergo an infinitesimal anticlockwise rotation, as shown in figure 5. We seem to have a propagator whose mass takes the wrong sign, i.e., a tachyon, which contributes a residue with an additional factor of ii. This tachyonic mode turns out to be spurious — a fact corroborated by our considerations in appendix A and continuity with the exterior contribution — but prevents us from calculating $G_{-}^{(2)}$ in closed form, at least by any method of which we are aware. Consequently, the renormalisation will have to be performed numerically. We elaborate on this in the next section.
@@ 5.2 Renormalisation
For a constant field $\varphi$, the one-loop contribution to the equation of motion is contained within the first derivative of the Coleman–Weinberg effective potential [27], given by
$V_{\text{CW}}(\phi)=V(\phi)+\frac{1}{2}\delta m^{2}\phi^{2}+\frac{1}{4}\delta\lambda\phi^{4}-\frac{i}{2}\int\,\frac{\text{d}^{4}p}{\left(2\pi\right)^{4}}\,\ln\left(1-\frac{1}{p^{2}}\frac{\text{d}^{2}V}{\text{d}\phi^{2}}\right)\>,$ (5.10)
@@
$\int\,\frac{\text{d}^{4}p}{\left(2\pi\right)^{4}}\,\ln\left(1-\frac{1}{p^{2}}\frac{\text{d}^{2}V}{\text{d}\phi^{2}}\right)$ $=\frac{i}{2\pi^{2}}\int_{0}^{\Lambda}\text{d}p\,p^{2}\left(\sqrt{p^{2}+\frac{\text{d}^{2}V}{\text{d}\phi^{2}}}-p\right)\>,$
assuming spherical symmetry. Imposing the renormalisation conditions [36],
@@
Since we are unable to determine the UV-divergent part of the interior Green’s function in closed form, we will need to integrate the Green’s function numerically. We do this using standard functions in Mathematica. To effect the renormalisation, we introduce pseudo-counterterms that allow us to subtract the divergences at the level of the integrand [36]. These are functions $\Delta m^{2}(p)$ and $\Delta\lambda(p)$ that satisfy the integral equations
@@
since it is terms like these that give rise to logarithmic divergences, but not quadratic divergences. Integrating and matching like terms fixes the constants to be
@@
Figure 6: The relative error $\eta$ between the exact and numerical calculations of the integral of $G_{+}^{(1)}$. Note that the horizontal axis begins at $\phi_{0}/v$, where $\phi_{0}$ is the value of the field at the surface of the source.
@@ 5.3 One-loop correction
With the tadpole contribution successfully renormalised, we can now solve the equation of motion for the quantum correction $\delta\phi$. The renormalised correction to the classical field configuration $\delta\phi$ is time dependent in general and satisfies
@@
Notice that the differential operator is that which defines the Green’s function for $k=E$, equivalently $n=2$, where $n$ is the order of the associated Legendre function. Therefore, this equation is solved by the convolution
@@
For the parameters used to generate figure 7, this amounts to a shift of about $8.5\%$ in the vacuum expectation value, which is precisely what we observe in figure 8. The more subtle feature hiding in figure 7 is that the quantum field is not simply a rescaled version of the classical field. The derivative of the field profile, when normalised relative to its VEV, has also shifted. Figure 9 shows this. The mass roughly determines the rate with which the field interpolates between its two asymptotes. Thus, it follows that the shift in the derivative is due to the shift in the relationship between the Lagrangian mass parameter $\mu$ and the physical mass $m$. At tree level, the physical mass is just the previously derived mass of scalar fluctuations,
@@
Since all functions here are spherically symmetric, it is sufficient to consider only the radial component of the force hereafter. We do not attempt to determine the exact way in which the mass and VEV shifts contribute to $\delta\phi$. However, from eq. (5.33), one might naively assume that the VEV shift contributes “twice”, in a sense (since the field is squared), while the mass shift (which applies because of the derivative) contributes “once”. This very rough argument implies that the relative shift in the force or acceleration $\Delta F$ is of order
@@
From a naive perturbative expansion, one might thus expect the shift in the fifth force to scale as $6\lambda/\pi^{2}$, that is remarkably close to the scaling which find numerically. This expression suggests that the shift in the force due to the self-interaction is, to a good approximation, linear in the coupling — at least in the perturbative regime — despite the non-perturbative dependence of the VEV on the coupling. Figure 12 corroborates this claim. For the parameters used to generate figure 7, the naive calculation predicts a $30\%$ shift in the strength of the force, which is precisely what we observe, around a distance of one Compton wavelength, in figure 11(a). This approximation holds well throughout the parameter space; the values used to generate figure 11(b) would imply a correction of about $20\%$, which the plot agrees with. Naturally, the approximation breaks down very close to or far from the surface of the source, where one would expect quantum effects to become less relevant. Notice, however, that, for large masses, we observe strong deviations from tree-level even close to the source.
The quantum-corrected force is generally weaker than its classical counterpart, except after about four Compton wavelengths from the surface of the source, where the quantum force becomes stronger. While this strengthening represents a relatively large fractional change, in absolute terms, the difference is negligible. As one would expect, the shift in the force vanishes as $\lambda\rightarrow 0$. Note, however, that, for sufficiently small $\lambda$, quantum corrections due to interactions with Standard Model fields are expected to become the dominant contribution, which we leave for future work.
@@
It is worth considering how the magnitude of the correction to the force depends on the other free parameters. Since changes in the mass scale $M$ are degenerate with changes in the density $\rho_{0}$ of the source, the only remaining free parameter of interest is the field mass $\mu$. Our rough arguments from earlier in this subsection would suggest that changing the mass parameter should have next to no effect on the force. This is approximately true. We observe a slight negative correlation between $\mu$ and $\Delta F$, but one that amounts to a change of about $0.4\%$ in the magnitude of the relative correction to the force for all $\mu\in(0,1]\,$GeV (the widest mass range where our numerical methods yield trustworthy results) in figure 12(c), with a similar relationship shown in figure 12(d). Note that we have extrapolated the linear relationship, ignoring the apparent divergence for small masses. This divergence may be a numerical artefact, a consequence of underflow errors. Alternatively, this could be a signal of the breakdown of the thin-wall approximation, since for $\mu\rightarrow 0$ we have $1/\mu\gg R$.
@@
[…]
@@ 6 Conclusions
We set out to estimate the magnitude of the leading-order quantum corrections to the symmetron field around extended sources. To this end, we have solved for the exact field profile around spherical sources whose radii are much larger than the Compton wavelength of the field, meaning our results are also relevant for systems with planar geometry [60, 59, 13, 14, 29, 65, 64]. We have computed the inverse of the fluctuation operator (the Green’s function) and the renormalised tadpole contribution. With that, we derived and numerically solved the equation of motion satisfied by the one-loop correction to the classical field profile. We have found that the action of quantum corrections is to flatten the gradients of the classical field profiles and reduce their vacuum expectation values, which agrees with similar analyses of the impact of quantum corrections on tunnelling configurations [36]. As a result, forces derived from one-loop field profiles are generally weaker than those computed from tree-level profiles, for the same input parameters.
The aim of this work was to obtain an analytical estimate of the relevance of quantum corrections to calculations of symmetron fifth forces. This is to enable a first quantification of the theoretical uncertainty on fifth-force calculations that results from ignoring radiative effects, as is commonly done in the literature. To make the problem analytically tractable, we employed the thin-wall and planar approximations, allowing us to treat spherical and planar geometries on the same footing. Of course, the variation of the fifth force far away from the surface of the source depends strongly on the dimensionality of the system, whether spherical, cylindrical or planar. However, as we have seen, the quantum corrections are maximal over a finite range close to the surface of the source where the gradients in the field profile and the nonlinearities in the field equations are largest. We therefore do not expect the dimensionality of the problem to have a significant impact on the percentage shift in the field due to quantum corrections. Quantifying the systematic error in this shift resulting from these approximations in the spherical case would require a numerical analysis of both the classical field profile and the resulting Green’s function. Such an analysis is beyond the scope of this work and may be presented elsewhere.
@@
In certain areas of parameter space, the classical field remains virtually identical to the quantum field. This is true in particular for very small self-coupling, as one might expect. One should note, however, that for $\lambda\lesssim 10^{-40}$, it is expected that couplings to the Standard Model dominate, and thus an additional source of quantum corrections would have to be taken into account. For sufficiently high but still perturbative couplings, we observe that the quantum-corrected force can be considerably weaker than the classical prediction. Indeed, for regions of parameter space that sit within current constraints [9, 67, 35], the symmetron force is as much as 30% weaker than classically predicted, for the same input parameters.
@@
While our results speak mostly to perturbative self-couplings, there is a clear trend that shows that the strength of the correction scales almost linearly with the strength of the self-interaction. We suspect that this trend continues into the non-perturbative regime and may even amount to a symmetron force that is practically undetectable for sufficiently high self-couplings. Phenomenologically, it would also be difficult to use such a field to account for observations related to dark matter, since even environments with low ambient density may have vanishingly small fifth forces. Our results also apply directly to systems with large, screened and spherical sources or systems with planar geometry. However, such an idealisation does not mean that we should expect quantum corrections to be minor for systems with, say, cylindrical symmetry [53], or relatively small sources. Field profiles still flatten away from the thin-wall approximation [39].
@@
We emphasise two different conclusions to draw from our work. The first is that, due to renormalisation, the relationship between Lagrangian parameters and physical observables shifts from the tree-level one. This conclusion alone would imply that current constraints apply to slightly different regions of parameter space. The second conclusion is that the spatial variation of the fifth force also changes, as indicated by figure 11, generally growing more slowly and peaking at a different point in space in the one-loop approximation. This could impact how we optimise geometries of future tabletop experiments [59]. The spatial variation of the quantum correction is a crucial point. Since $\delta\phi$ changes with position, there is no point in space that we could choose to define the physical mass and self-coupling such that the quantum correction vanishes everywhere. Put differently, there does not exist a convenient choice of renormalisation scheme that makes the quantum corrections disappear. They cannot be fine-tuned away, as it were. The spatially varying quantum fluctuations about the classical background are an intrinsic and unavoidable feature of the theory.
@@
We propose several directions for future work. First, the size of the one-loop corrections suggests that higher-order corrections may still have measurable contributions. To help with this endeavour, by use of the Schwinger-Keldysh closed-time path formalism [58, 47], it may be possible to simplify the portion of our analysis concerned with deriving an expression for quantum-corrected observables and the quantum field, and perhaps even obtain its equation of motion, all without reference to the effective action. This may be presented in future work. Second, the method presented in this paper can likely produce estimates of quantum corrections for symmetron-like models, such as the complex symmetron [3] or generalised symmetron [45]. Third, it has been shown that one can get around some experimental constraints by choosing models in which screening arises at the one-loop level [19], provided the background field is constant [37]. The methods developed in this paper would allow us to revisit this result and take into account the spatial variation of the field profile. In addition, a future investigation could consider the effect of quantum corrections for a broader range of observables, including shifts in frequency spectra [46], properties of white dwarfs (mass, radius, luminosity, etc.) [3] and gravitational lensing [45]. Finally, we hope that this or a similar analysis could be used to further explore the quantum nature of fifth-force theories in general.